# Stronger Constraints on Primordial Black Holes as Dark Matter Derived from the Thermal Evolution of the Intergalactic Medium over the Last Twelve Billion Years

### v1

@@ IV Modifications to DarkHistory
In order to solve our thermochemistry, we have modified the publicly available code DarkHistory [76]. We make three modifications to this code: (a) addition of PBH energy injection, (b) inclusion of He iii evolution, and (c) extension of the calculations to $z=0$. We now describe these modifications. In order to incorporate energy injection by PBHs, we include the electron and positron spectra from Equation (3) a single PBH into the in_spec_elec data structure and the photon spectrum, again following from Equation (3), into the in_spec_phot data structure in the main.evolve function. Additionally, we modify the phys.inj_rate function to account for the total energy injection rate from PBHs according to Equation (4). To include He iii evolution, we modify the function tla.get_history, which solves the temperature and ionization evolution equations at each time step $\Delta t$. Instead of setting $x_{\mathrm{HeIII}}$ to zero, as is the case in the public version of DarkHistory, we now self-consistently use the values resulting from an equation of the helium ionization equations at each time step. Finally, to extend the calculations to $z=0$, we modify the main.evolve function. Incorporating He iii chemistry and achieving calculations down to $z=0$ also necessitates recalculating the deposition fractions $f_{\mathrm{Cnl}}$. In the public version of DarkHistory, values of $f_{\mathrm{Cnl}}$ are precomputed for a range of redshifts and ionization fractions, without considering He iii chemistry and only extending down to $z\sim 3$. To overcome these constraints, we adapt DarkHistory to calculate $f_{\mathrm{Cnl}}$ dynamically and self-consistently at each time step. This is accomplished by integrating the deposition fraction computation directly into the main.evolve function, simultaneously incorporating He iii chemistry. This enhancement in calculating deposition fractions represents the central methodological improvement in this paper compared to previous works. We initialise DarkHistory at $z=3,000$ using default initial values for all quantities.
@@ V Results from Modified DarkHistory
Using the coupled Equations 8, 10 and 13, we calculate the temperature and ionization histories as functions of redshift for a range of $M_{\mathrm{PBH}}$ and $f_{\mathrm{PBH}}$ values. These calculations begin at redshift $30001+z=$ 3000 with initial conditions $x_{\mathrm{HII}}=1$ and $x_{\mathrm{HeII}}=1$. The model-predicted temperatures are then compared with observational data, which will be described in detail in the next section, for the redshift range $5.80.06<z<$ 5.8. Figure 3 shows the temperature evolution calculated for $M_{\mathrm{PBH}}=7\times 10^{16}$ g, along with the dataset. We examine $f_{\mathrm{PBH}}$ values of 0, $5\times 10^{-3}$, and $5\times 10^{-2}$. A value of 0 indicates the absence of PBHs, and this baseline model fits the data very well. For the other two $f_{\mathrm{PBH}}$ values, the temperature increases at lower redshifts due to enhanced Coulomb heating (as seen in Figure 2), with this effect being most prominent at the lowest redshifts. By visual inspection, we can exclude the $f_{\mathrm{PBH}}=5\times 10^{-2}$ model, as it significantly overpredicts the IGM temperature and does not provide a good fit to the data. The evolution of the ionization fraction is shown in the right panel of Figure 3, where the ionization fraction remains unchanged after the inclusion of PBHs. This is expected, as the $f_{\mathrm{ion}}$ values for ionization are close to zero, even for large $f_{\mathrm{PBH}}$ values such as $5\times 10^{-2}$. This suggests that PBH-induced energy injection predominantly affects the thermal rather than the ionization properties of the IGM.
@@
Figure 3 also demonstrates how our analysis improves upon Ref. [40], enabling more stringent constraints on PBH dark matter from IGM temperature measurements. A critical distinction between our approach and Ref. [40] is their assumption that $x_{\mathrm{HeIII}}=0$. However, neglecting He iii evolution impacts the ionization state of the IGM, x, and the free electron number density, $n_{e}$. This, in turn, affects the energy deposition efficiency, $f_{\mathrm{heat}}(z,\textbf{x})$ (Equation 7), and consequently modifies the heating rate, $\dot{T}_{\mathrm{PBH}}$ (Equation 9). Additionally, omitting He iii alters both photoionization and photoheating rates, resulting in an unphysical temperature evolution.
@@
To explicitly evaluate the impact of neglecting He iii, we performed a comparative calculation by artificially setting $x_{\mathrm{HeIII}}=0$ and contrasting it with our comprehensive model, which fully incorporates He iii evolution. As illustrated in Figure 3, the gray curve corresponds to the model without He iii, while the orange curve represents our full calculation for $M_{\mathrm{PBH}}=7\times 10^{16}$g and $f_{\mathrm{PBH}}=5\times 10^{-3}$. This comparison clearly demonstrates that neglecting He iii initiates a significant deviation in temperature at, in this case, $z\lesssim 7$ (left panel), which intensifies towards lower redshifts. Furthermore, omitting He iii evolution noticeably impacts the He ii fraction beginning around $z\sim 7$ (right panel). Intriguingly, Ref. [40] also reported a similar, unexpected temperature increase, visible as a teal curve in Figure 1 of their paper. Our findings strongly suggest that this temperature anomaly results directly from neglecting He iii evolution, emphasizing the necessity of including comprehensive helium evolution to accurately constrain PBH abundance.
@@
Furthermore, the analysis in Ref. [40] disregards IGM temperature data at $z<3.5$, thereby significantly weakening their constraints. As Figure 3 demonstrates, cumulative heat injection from PBHs becomes particularly substantial at lower redshifts, highlighting the importance of extending the analysis to $z=0$.
@@
A third critical distinction pertains to the treatment of reionization and photoheating. Ref. [40] employs a simplified reionization model that does not self-consistently address IGM thermochemistry or incorporate current understandings of hydrogen reionization and its subsequent impacts. In contrast, we adopt the physically motivated reionization model by Puchwein et al. [77], which explicitly models ionizing radiation from known populations of active galactic nuclei (AGN) and star-forming galaxies, together with a realistic density distribution of the inhomogeneous IGM. The Puchwein et al. model also compared favourably with more sophisticate three-dimensional radiative transfer cosmological simulations of the IGM that are known to reproduce several IGM observations. Conversely, Ref. [40] merely adjust their reionization history to align with Planck measurements without explicitly considering the astrophysical sources. Their “photoheated” model simplifies pre-reionization heating as $\dot{T}\star=\dot{x}_{\mathrm{HII}}(1+\chi)\Delta T$, where $\Delta T$ is an empirical parameter ranging between $2\times 10^{4}$ K and $3\times 10^{4}$ K, and post-reionization heating relies solely on the spectral index of the ionizing background. While flexible, this approach fails to explicitly track contributions from distinct astrophysical mechanisms.
@@
In summary, our analysis differs crucially from Ref. [40] by explicitly accounting for He iii evolution, extending the analysis down to redshift zero, and adopting a physically comprehensive reionization model. Overall, on the one hand, these improvements reduce the temperature enhancement due to PBHs. On the other hand, these changes allow us to use measurements from a much wider redshift range than before, which leads directly to significantly tighter constraints on the abundance of PBH dark matter, as will be discussed further below.
@@ VI The Temperature Measurements of the Intergalactic Medium from the Lyman-$\alpha$ Forest
[…]
@@ VII Constraints on PBHs from IGM temperature
Figure 4: Our limits on $f_{\mathrm{PBH}}$ for PBH masses between $10^{15}\,\mathrm{g}$ and $2\times 10^{17}\,\mathrm{g}$, inferred at the 95% confidence level from measurements of the IGM temperature between redshifts 0 and 6, are shown in the figure. The shaded region represents the excluded parameter space based on observational data. For comparison, we also show similar constraints from Ref. [40], depicted in gray, where we have adopted their strongest reported constraint from the “photoheating” model variations. Our constraints are stronger than those from Ref. [40] by an order of magnitude or more.
@@
The resulting constraints on PBH parameters are shown in Figure 4. The red curve represents the upper limit on $f_{\mathrm{PBH}}$ as a function of PBH mass, $M_{\mathrm{PBH}}$. The red-shaded region corresponds to the excluded parameter space at the 95% confidence level. At $M_{\mathrm{PBH}}=7\times 10^{16}$ g, we constrain $f_{\mathrm{PBH}}\leq 8.6\times 10^{-3}$. The constraints weaken for larger PBH masses due to the slower evaporation rate of more massive PBHs. Since Hawking radiation causes PBHs to lose mass at a rate $\dot{M}_{\mathrm{PBH}}\propto M_{\mathrm{PBH}}^{-2}$ [79, 80], the total energy injection rate per unit volume, given by $(\mathrm{d}E/\mathrm{d}V\mathrm{d}t)_{\mathrm{inj}}\propto n_{\mathrm{PBH}}\dot{M}_{\mathrm{PBH}}$, scales as $M_{\mathrm{PBH}}^{-3}$ as the PBH number density follows $n_{\mathrm{PBH}}\propto M_{\mathrm{PBH}}^{-1}$. As a result, energy injection decreases with increasing $M_{\mathrm{PBH}}$, leading to weaker constraints.
@@
For comparison, we show previous constraints based on the IGM temperature, from Ref. [40] by the gray curve in Figure 4, where we have adopted their strongest reported constraint from the ‘photoheating’ model variations. Our constraints are stronger than these across all mass ranges, with the improvement being more than an order of magnitude for some PBH mass values. As we discussed above, this improvement arises from our inclusion of low-redshift temperature data, where PBH-induced heating via Coulomb interactions is most significant. Additionally, our constraints are also different due to the incorporation of He iii evolution and a more realistic reionization model [77].
@@
Relative to other constraints in the literature (not shown in Figure 4, in order to keep the figure accessible), our constraints are the second strongest in the mass range $7\times 10^{15}$ g to $2\times 10^{17}$ g [41, 42, 44, 45, 46, 47, 49, 50]. The most stringent limit in this mass window currently comes from measurements of the 511 keV gamma-ray line [54]. While this provides a stronger constraint, it crucially depends on astrophysical modeling of positron propagation in the Galactic magnetic fields, and dark matter density profiles, introducing significant systematic uncertainties. In comparison, the Lyman-$\alpha$ forest has been extensively used to derive a wide range of cosmological constraints, demonstrating its precision and robustness as a probe of the IGM. Our IGM temperature constraints, based on well-understood cosmological heating and cooling processes, are independent of local PBH density and astrophysical backgrounds. Additionally, while the 511 keV gamma-ray signal relies on indirect annihilation signatures, IGM temperature measurements offer a more direct window into PBH-induced heating. Given these distinctions, the Lyman-$\alpha$ forest may provide a complementary and potentially more reliable avenue for constraining PBHs.
@@
Note that, in Figure 4, we consider PBHs with masses greater than $\sim 10^{15}$ g, as those below this threshold would have evaporated by now, given their lifetimes are shorter than the age of the universe [69, 70, 71, 72]. This sets the lower limit for PBH masses in our analysis. However, see recent studies [81, 82, 83, 84] for discussions on how quantum-gravity effects might still allow PBHs with $M_{\rm PBH}\ll 10^{15}$ g to exist today.

### v2

@@ IV Modifications to DarkHistory
In order to solve our thermochemistry, we have modified the publicly available code DarkHistory [2020PhRvD.101b3530L]. We make three modifications to this code: (a) addition of PBH energy injection, (b) inclusion of He iii evolution, and (c) extension of the calculations to $z=0$. We now describe these modifications. In order to incorporate energy injection by PBHs, we include the electron and positron spectra from Equation (3) a single PBH into the in_spec_elec data structure and the photon spectrum, again following from Equation (3), into the in_spec_phot data structure in the main.evolve function. Additionally, we modify the phys.inj_rate function to account for the total energy injection rate from PBHs according to Equation (4). To include He iii evolution, we modify the function tla.get_history, which solves the temperature and ionization evolution equations at each time step $\Delta t$. Instead of setting $x_{\mathrm{HeIII}}$ to zero, as is the case in the public version of DarkHistory, we now self-consistently use the values resulting from an equation of the helium ionization equations at each time step. Finally, to extend the calculations to $z=0$, we modify the main.evolve function. Incorporating He iii chemistry and achieving calculations down to $z=0$ also necessitates recalculating the deposition fractions $f_{\mathrm{Cnl}}$. In the public version of DarkHistory, values of $f_{\mathrm{Cnl}}$ are precomputed for a range of redshifts and ionization fractions, without considering He iii chemistry and only extending down to $z\sim 3$. To overcome these constraints, we adapt DarkHistory to calculate $f_{\mathrm{Cnl}}$ dynamically and self-consistently at each time step. This is accomplished by integrating the deposition fraction computation directly into the main.evolve function, simultaneously incorporating He iii chemistry. This enhancement in calculating deposition fractions represents the central methodological improvement in this paper compared to previous works. We initialise DarkHistory at $z=3,000$ using default initial values for all quantities.
Although the original DarkHistory code was developed for high-redshift applications ($z\gtrsim 3$) and primarily models hydrogen and singly ionized helium, our modified version extends its validity down to $z=0$ by including He iii evolution and adopting the reionization model of Puchwein et al. [2019MNRAS.485...47P]. This extension allows a self-consistent treatment of late-time photoheating and photoionization. Since PBH evaporation acts as a decay process, with energy injection scaling linearly with PBH density, it is not enhanced by clustering or halo formation, making our constraints robust to clustering uncertainties. However, we note that spatial inhomogeneities and detailed astrophysical feedback from galaxies and AGNs are not explicitly modeled. The use of mean-density photoheating rates ensures realistic average thermal histories, but small-scale variations and local reionization effects remain beyond the scope of this study. Future work could incorporate these aspects through full radiative-transfer simulations.
@@ V Results from Modified DarkHistory
Using the coupled Equations 8, 10 and 13, we calculate the temperature and ionization histories as functions of redshift for a range of $M_{\mathrm{PBH}}$ and $f_{\mathrm{PBH}}$ values. These calculations begin at redshift $1+z=3000$ with initial conditions $x_{\mathrm{HII}}=1$ and $x_{\mathrm{HeII}}=1$. The model-predicted temperatures are then compared with observational data, which will be described in detail in the next section, for the redshift range $0.06<z<5.8$. Figure 3 shows the temperature evolution calculated for $M_{\mathrm{PBH}}=7\times 10^{16}$ g, along with the dataset. We examine $f_{\mathrm{PBH}}$ values of 0, $5\times 10^{-3}$, and $5\times 10^{-2}$. A value of 0 indicates the absence of PBHs, and this baseline model fits the data very well. For the other two $f_{\mathrm{PBH}}$ values, the temperature increases at lower redshifts due to enhanced Coulomb heating (as seen in Figure 2), with this effect being most prominent at the lowest redshifts. By visual inspection, we can exclude the $f_{\mathrm{PBH}}=5\times 10^{-2}$ model, as it significantly overpredicts the IGM temperature and does not provide a good fit to the data. The evolution of the ionization fraction is shown in the right panel of Figure 3, where the ionization fraction remains unchanged after the inclusion of PBHs, with differences always below 1%. This is expected, as the $f_{\mathrm{ion}}$ values for ionization are close to zero, even for large $f_{\mathrm{PBH}}$ values such as $5\times 10^{-2}$. This suggests that PBH-induced energy injection predominantly affects the thermal rather than the ionization properties of the IGM.
@@
Figure 3 also demonstrates how our analysis improves upon Ref. [2024arXiv240910617S], enabling more stringent constraints on PBH dark matter from IGM temperature measurements. A critical distinction between our approach and Ref. [2024arXiv240910617S] is their assumption that $x_{\mathrm{HeIII}}=0$. However, neglecting He iii evolution impacts the ionization state of the IGM, x, and the free electron number density, $n_{e}$. This, in turn, affects the energy deposition efficiency, $f_{\mathrm{heat}}(z,\textbf{x})$ (Equation 7), and consequently modifies the heating rate, $\dot{T}_{\mathrm{PBH}}$ (Equation 9). Additionally, omitting He iii alters both photoionization and photoheating rates, resulting in an unphysical temperature evolution.
@@
To explicitly evaluate the impact of neglecting He iii, we performed a comparative calculation by artificially setting $x_{\mathrm{HeIII}}=0$ and contrasting it with our comprehensive model, which fully incorporates He iii evolution. As illustrated in Figure 3, the gray curve corresponds to the model without He iii, while the orange curve represents our full calculation for $M_{\mathrm{PBH}}=7\times 10^{16}$g and $f_{\mathrm{PBH}}=5\times 10^{-3}$. This comparison clearly demonstrates that neglecting He iii initiates a significant deviation in temperature at, in this case, $z\lesssim 7$ (left panel), which intensifies toward lower redshifts. Furthermore, omitting He iii evolution noticeably impacts the He ii fraction beginning around $z\sim 7$ (right panel). Intriguingly, Ref. [2024arXiv240910617S] also reported a similar, unexpected temperature increase, visible as a teal curve in Figure 1 of their paper. Our findings strongly suggest that this temperature anomaly results directly from neglecting He iii evolution, emphasizing the necessity of including comprehensive helium evolution to accurately constrain PBH abundance.
@@
Furthermore, the analysis in Ref. [2024arXiv240910617S] disregards IGM temperature data at $z<3.5$, thereby significantly weakening their constraints. As Figure 3 demonstrates, cumulative heat injection from PBHs becomes particularly substantial at lower redshifts, highlighting the importance of extending the analysis to $z=0$.
@@
A third critical distinction pertains to the treatment of reionization and photoheating. Ref. [2024arXiv240910617S] employs a simplified reionization model that does not self-consistently address IGM thermochemistry or incorporate current understandings of hydrogen reionization and its subsequent impacts. In contrast, we adopt the physically motivated reionization model by Puchwein et al. [2019MNRAS.485...47P], which explicitly models ionizing radiation from known populations of active galactic nuclei (AGN) and star-forming galaxies, together with a realistic density distribution of the inhomogeneous IGM. The Puchwein et al. model also compared favorably with more sophisticate three-dimensional radiative transfer cosmological simulations of the IGM that are known to reproduce several IGM observations. Conversely, Ref. [2024arXiv240910617S] merely adjust their reionization history to align with Planck measurements without explicitly considering the astrophysical sources. Their “photoheated” model simplifies pre-reionization heating as $\dot{T}\star=\dot{x}_{\mathrm{HII}}(1+\chi)\Delta T$, where $\Delta T$ is an empirical parameter ranging between $2\times 10^{4}$ K and $3\times 10^{4}$ K, and post-reionization heating relies solely on the spectral index of the ionizing background. While flexible, this approach fails to explicitly track contributions from distinct astrophysical mechanisms.
@@
In summary, our analysis differs crucially from Ref. [2024arXiv240910617S] by explicitly accounting for He iii evolution, extending the analysis down to redshift zero, and adopting a physically comprehensive reionization model. Overall, on the one hand, these improvements reduce the temperature enhancement due to PBHs. On the other hand, these changes allow us to use measurements from a much wider redshift range than before, which leads directly to significantly tighter constraints on the abundance of PBH dark matter, as will be discussed further below.
@@ VI The Temperature Measurements of the Intergalactic Medium from the Lyman-$\alpha$ Forest
We now describe the IGM temperature measurements presented in Figure 3 and used below to derive constraints on PBHs. Gaikwad et al. 2020 [2020MNRAS.494.5091G] investigated high-resolution Lyman-$\alpha$ spectra from QSOs at $6<z\lesssim 6.5$ to study the incidence of Lyman-$\alpha$ transmission spikes. By comparing the incidence of Lyman-$\alpha$ transmission spikes with IGM simulations, they found that the width and height of the transmission spikes are sensitive to the IGM’s temperature. This lead to an inference of $T_{\mathrm{IGM}}=(11,000\pm 1600,10,500\pm 2100,12,000\pm 2200)$ K at $z=5.4,5.6,5.8$ respectively. Boera et al. 2019 [2019ApJ...872..101B] use a power-law model for the IGM temperature density relation to measure the thermal state of the IGM at $4<z<5.2$ using 1D Lyman-$\alpha$ transmission power spectrum data from 15 quasar spectra obtained from Keck/HIRES and VLT/UVES. They also introduced a new parameter, $u_{0}$, to account for pressure smoothing, representing the cumulative energy per proton deposited into a gas parcel at the mean background density. From their analysis, they inferred a temperature $T_{\mathrm{IGM}}\sim 7000$80008000 K with no significant evolution with redshift. In a separate study, Gaikwad et al. 2021 [2021MNRAS.506.4389G] focused on the thermal state of the IGM in the redshift range $2\leq z\leq 4$. They analyzed 103 high-resolution Lyman-$\alpha$ spectra using four transmission distribution statistics: the transmission power spectrum, Doppler parameter distribution, wavelet statistics, and curvature statistics. By comparing these measurements with cosmological hydrodynamic simulations, they inferred a temperature of $T_{\mathrm{IGM}}=14,750\pm 1,322$ K at $z\sim 3$, which they attributed to He ii reionization. Recently, Hu et al. 2025 [2025MNRAS.536....1H] extended IGM temperature measurements to lower redshifts, providing the first direct constraints on the thermal and ionization state of the IGM at $0.9<z<1.5$ by modeling the distribution of line widths and neutral hydrogen column density values of individual absorption lines in the Lyman-$\alpha$ forest. They analyzed 301 Lyman-$\alpha$ absorption lines from 12 archival HST spectra. Their measurements yielded the following IGM temperatures at mean density: $T_{\mathrm{IGM}}=13,520\,\mathrm{K}$ at $z=1.4$, $T_{\mathrm{IGM}}=6,020\,\mathrm{K}$ at $z=1.2$, and $T_{\mathrm{IGM}}=13,080\,\mathrm{K}$ at $z=1$. Ricotti et al. 2000 [2000ApJ...534...41R] measured the temperature parameter at very low redshift using the line width and column density distribution. They analyzed 79 lines from 15 Seyfert galaxies and quasars detected in observations with the HST/GHRS. Their results yielded a temperature of $T_{\mathrm{IGM}}=4,700$ K at $z=0.06$, with an error of $1,739$ K. All of these data are shown in the left panel of Figure 3. We now use these measurements to derive constraints on PBHs.
@@ VII Constraints on PBHs from IGM temperature
Figure 4: Our limits on $f_{\mathrm{PBH}}$ for PBH masses between $10^{15}\,\mathrm{g}$ and $2\times 10^{17}\,\mathrm{g}$, inferred at the 95% confidence level from measurements of the IGM temperature between redshifts 0 and 6, are shown in the figure. The shaded region represents the excluded parameter space based on observational data. For comparison, we also show similar constraints from Ref. [2024arXiv240910617S], depicted in gray, where we have adopted their strongest reported constraint from the “photoheating” model variations. Our constraints are stronger than those from Ref. [2024arXiv240910617S] by an order of magnitude or more.
@@
The resulting constraints on PBH parameters are shown in Figure 4. The red curve represents the upper limit on $f_{\mathrm{PBH}}$ as a function of PBH mass, $M_{\mathrm{PBH}}$. The red-shaded region corresponds to the excluded parameter space at the 95% confidence level. At $M_{\mathrm{PBH}}=7\times 10^{16}$ g, we constrain $f_{\mathrm{PBH}}\leq 8.6\times 10^{-3}$. The constraints weaken for larger PBH masses due to the slower evaporation rate of more massive PBHs. Since Hawking radiation causes PBHs to lose mass at a rate $\dot{M}_{\mathrm{PBH}}\propto M_{\mathrm{PBH}}^{-2}$ [1974Natur.248...30H, MacGibbon:1991tj], the total energy injection rate per unit volume, given by $(\mathrm{d}E/\mathrm{d}V\mathrm{d}t)_{\mathrm{inj}}\propto n_{\mathrm{PBH}}\dot{M}_{\mathrm{PBH}}$, scales as $M_{\mathrm{PBH}}^{-3}$ as the PBH number density follows $n_{\mathrm{PBH}}\propto M_{\mathrm{PBH}}^{-1}$. As a result, energy injection decreases with increasing $M_{\mathrm{PBH}}$, leading to weaker constraints. Here, we note that the constraint on PBH parameters assumes a monochromatic mass distribution. For an extended mass distribution, such as the commonly used log-normal mass function [Dasgupta:2019cae], the constraint would extend over a broader mass range, probing more into the asteroid-mass range $(10^{17}\,\textrm{g}\leq M_{\rm PBH}\leq 10^{22}\,\textrm{g}$).
@@
For comparison, we show previous constraints based on the IGM temperature, from Ref. [2024arXiv240910617S] by the gray curve in Figure 4, where we have adopted their strongest reported constraint from the ‘photoheating’ model variations. Our constraints are stronger than these across all mass ranges, with the improvement being more than an order of magnitude for some PBH mass values. As we discussed above, this improvement arises from our inclusion of low-redshift temperature data, where PBH-induced heating via Coulomb interactions is most significant. Additionally, our constraints are also different due to the incorporation of He iii evolution and a more realistic reionization model [2019MNRAS.485...47P].
@@
Relative to other constraints in the literature (not shown in Figure 4, in order to keep the figure accessible), our constraints are the second strongest in the mass range $7\times 10^{15}$ g to $2\times 10^{17}$ g [Carr:2009jm, Boudaud:2018hqb, DeRocco:2019fjq, Laha:2019ssq, Dasgupta:2019cae, Laha:2020ivk, Ray:2021mxu, Coogan:2020tuf] (a comprehensive comparison with existing bounds is presented in Appendix D). The most stringent limit in this mass window currently comes from measurements of the 511 keV gamma-ray line [2024PhRvD.110l3022L]. While this provides a stronger constraint, it crucially depends on astrophysical modeling of positron propagation in the Galactic magnetic fields, and dark matter density profiles, introducing significant systematic uncertainties (see Ref. [DeRocco:2019fjq, Laha:2019ssq, Dasgupta:2019cae] for weaker constraints derived from the 511 keV measurement and discussions on how this constraint can relax under various astrophysical uncertainties). In comparison, the Lyman-$\alpha$ forest has been extensively used to derive a wide range of cosmological constraints, demonstrating its precision and robustness as a probe of the IGM. Our IGM temperature constraints, based on well-understood cosmological heating and cooling processes, are independent of local dark matter density and astrophysical backgrounds. Additionally, while the 511 keV gamma-ray signal relies on indirect annihilation signatures, IGM temperature measurements offer a more direct window into PBH-induced heating. Given these distinctions, the Lyman-$\alpha$ forest may provide a complementary and potentially more reliable avenue for constraining PBHs.
@@
Note that, in Figure 4, we consider PBHs with masses greater than $\sim 10^{15}$ g, as those below this threshold would have evaporated by now, given their lifetimes are shorter than the age of the Universe [Page:1976df, Page:1976ki, MacGibbon:1990zk, MacGibbon:2007yq]. This sets the lower limit for PBH masses in our analysis. However, see recent studies [Dvali:2020wft, Thoss:2024hsr, Dvali:2024hsb, Basumatary:2024uwo] for discussions on how quantum-gravity effects might still allow PBHs with $M_{\rm PBH}\ll 10^{15}$ g to exist today.